### v1

# Minimax optimal differentially private synthetic data for smooth queries

Abstract
Differentially private synthetic data enables the sharing and analysis of sensitive datasets while providing rigorous privacy guarantees for individual contributors. A central challenge is to achieve strong utility guarantees for meaningful downstream analysis. Many existing methods ensure uniform accuracy over broad query classes, such as all Lipschitz functions, but this level of generality often leads to suboptimal rates for statistics of practical interest. Since many common data analysis queries exhibit smoothness beyond what worst-case Lipschitz bounds capture, we ask whether exploiting this additional structure can yield improved utility.
We study the problem of generating $(\varepsilon,\delta)$-differentially private synthetic data from a dataset of size $n$ supported on the hypercube $1,1]^{d}$, with utility guarantees uniformly for all smooth queries having bounded derivatives up to order $k$. We propose a polynomial-time algorithm that achieves a minimax error rate of $n^{-\min\{1,\frac{k}{d}\}}$, up to a $\log(n)$ factor. This characterization uncovers a phase transition at $k=d$. Our results generalize the Chebyshev moment matching framework of [48, 60] and strictly improve the error rates for $k$-smooth queries established in [60]. Moreover, we establish the first minimax lower bound for the utility of $(\varepsilon,\delta)$-differentially private synthetic data with respect to $k$-smooth queries, extending the Wasserstein lower bound for $\varepsilon$-differential privacy in [11].
1 Introduction
Differential privacy (DP) has emerged as a leading standard for safeguarding privacy in settings that require statistical analysis over large and sensitive datasets. Introduced to provide a rigorous, attack-agnostic notion of privacy protection, differential privacy limits what an adversary can infer about any single individual, even when the adversary has substantial auxiliary information [26]. Formally, a randomized mechanism is differentially private if its output distributions are nearly indistinguishable on any pair of adjacent datasets that differ in only one individual’s record. This framework is increasingly adopted in practice, most prominently in the disclosure avoidance system for the 2020 U.S. Census [2, 35, 34], and in large-scale deployments by technology companies [22, 25]. Beyond its foundational appeal, differential privacy has enabled a broad range of data science workflows, including private query answering and workload optimization [46], regression and empirical risk minimization [21, 53], parameter estimation [24], and private stochastic gradient methods for modern machine learning [50, 1].
Much of the existing DP literature focuses on designing task-specific mechanisms tailored to a predefined workload (e.g., a fixed set of queries or a particular model class). While often highly accurate, these approaches can require substantial expertise, ranging from sensitivity analysis and mechanism selection (e.g., Laplace/Gaussian noise, exponential mechanism [26]) to careful privacy accounting across iterative computations. A complementary and increasingly practical alternative is differentially private synthetic data generation, where one releases a synthetic dataset that approximates statistical properties of the original data while satisfying differential privacy [32, 9, 61, 8]. By the post-processing property [26] of differential privacy, analysts can subsequently perform a wide range of downstream tasks on the released synthetic dataset without incurring additional privacy loss.
1.1 Differentially private synthetic data
Differentially private synthetic data consist of artificially generated records that aim to preserve the aggregate statistical structure of a sensitive source dataset (often user-generated), while providing rigorous privacy guarantees to the individuals who contributed to the data [49]. By decoupling data utility from direct access to raw records, DP synthetic data offer a principled way to unlock the value of datasets that are otherwise difficult to share due to privacy concerns.
The task of generating private synthetic data can be formulated as follows. Let $(\Omega,\rho)$ be a metric space and consider a dataset $X=(X_{1},\dots,X_{n})\in\Omega^{n}$. The goal is to design an efficient randomized algorithm that outputs a differentially private synthetic dataset $Y=(Y_{1},\dots,Y_{m})\in\Omega^{m}$ such that the empirical measures
$p_{X}=\frac{1}{n}\sum_{i=1}^{n}\delta_{X_{i}}\quad\text{and}\quad p_{Y}=\frac{1}{m}\sum_{j=1}^{m}\delta_{Y_{j}}$
are close under an appropriate notion of discrepancy. A large body of prior work evaluates the utility of DP synthetic data by requiring accuracy for a prescribed finite collection of queries [33, 32, 15, 14, 30]. At the same time, there are fundamental computational barriers: under standard cryptographic assumptions, [55] showed that generating differentially private synthetic Boolean data that are simultaneously accurate for rich classes of Boolean queries is computationally intractable in general.
Motivated by the desire for task-agnostic utility guarantees, a more recent line of work measures the quality of synthetic data in a continuous space $\Omega$ via $\mathbb{E}\,W_{1}(p_{X},p_{Y})$, where $W_{1}$ denotes the $1$-Wasserstein distance and the expectation is taken over the randomness of the algorithm. By the Kantorovich–Rubinstein duality (see, e.g., [57]),
$W_{1}(p_{X},p_{Y})=\sup_{\mathrm{Lip}(f)\leq 1}\left(\int f\,\mathrm{d}p_{X}-\int f\,\mathrm{d}p_{Y}\right),$ (1)
where the supremum ranges over all $1$-Lipschitz functions on $\Omega$. Since many learning objectives and evaluation functionals are Lipschitz, or can be controlled by Lipschitz losses [59, 44, 13, 47], (1) yields a uniform performance guarantee for a broad range of downstream tasks performed on synthetic datasets whose empirical distribution is close to $p_{X}$ in $W_{1}$.
Within this Wasserstein framework, [11] gave the first $\varepsilon$-DP algorithm achieving the minimax rate $(\varepsilon n)^{-1/d}$ (up to polylogarithmic factors) for data in the hypercube $0,1]^{d}$ under the $\ell_{\infty}$ metric. Subsequently, [37] developed two classes of algorithms that attain the same minimax-optimal $(\varepsilon n)^{-1/d}$ rate for $d\geq 2$. These constructions have enabled the analysis of several downstream learning tasks [31, 62, 18]. Under the relaxed notion of $(\varepsilon,\delta)$-DP, [48] analyzed a Chebyshev moment matching algorithm that achieves the same $n^{-1/d}$ Wasserstein rate.
Recent work has sought to surpass the worst-case $n^{-1/d}$ scaling by exploiting additional structure in the data. For example, [36] considered datasets concentrated near a $d^{\prime}$-dimensional subspace and obtained a rate $(\varepsilon n)^{-1/d^{\prime}}$ up to polynomial factors in $d$, while [23] studied data supported on an unknown $d^{\prime}$-dimensional manifold in $0,1]^{d}$. When the dataset satisfies suitable sparsity conditions, [39] showed that the error rate can be further improved via an adaptation of the Private Measure Mechanism introduced in [37]. In a related direction, [29] analyzed the practical Private Evolution algorithm [45] through its connection to the Private Signed Measure Mechanism in [37]. Relatedly, [27] studied instance-optimal private density estimation under Wasserstein loss, and [38] extended these ideas to online synthetic data release.
Beyond Wasserstein distance, several alternative discrepancy measures have been considered, including $L_{1}$-type distances [61] and more general integral probability metrics (IPMs) [6]. The $1$-Wasserstein distance is a particular example of an IPM, and [12] characterized the sample complexity of DP synthetic data generation for unbounded query classes under IPMs. Moreover, [23] analyzed the class of $s$-sparse queries and obtained an $(\varepsilon n)^{-1/s}$ error rate.
While Wasserstein-based guarantees are appealing due to their uniform control over all $1$-Lipschitz test functions, they can be overly pessimistic for smoother statistics. In practice, many summaries of interest correspond to smooth queries with bounded derivatives, such as low-order moments and correlations $f({\mathbf{x}})=\prod_{j\in J\subset[d]}x_{j}$, linear statistics $f({\mathbf{x}})=\langle v,{\mathbf{x}}\rangle$ [10, 33], quadratic functionals $f(x)={\mathbf{x}}^{\top}A{\mathbf{x}}$ arising in covariance and correlation estimation [4], and smooth loss or score functions used in private regression, classification, and convex optimization [20, 56] such as Gaussian kernel function $f({\mathbf{x}})=\exp(\frac{\|{\mathbf{x}}-\mu\|^{2}}{2\sigma^{2}})$ and logistic function $f(x)=\frac{1}{1+e^{-x}}$. For such queries, utility guarantees tailored to smoothness can be substantially sharper than worst-case Lipschitz bounds [60]. This observation motivates the following question:
Can we generate DP synthetic data with utility guarantees for smooth queries that improve upon the $n^{-1/d}$ rate?
Prior work [60] studied the release of $k$-smooth queries under differential privacy and achieved an error rate of $O\left(n^{-\frac{2k}{3d+2k}}\right)$ (up to its dependence on the privacy parameter) under $(\varepsilon,\delta)$-DP and a rate $\left(n^{-\frac{k}{2d+k}}\right)$ under $\varepsilon$-DP. In the special case $k=1$, this rate does not recover the minimax-optimal $n^{-1/d}$ exponent established in [11]. In contrast, we improve the dependence on the sample size to $\tilde{O}\!\left(n^{-\min\{1,\frac{k}{d}\}}\right)$ and show that this rate is minimax optimal up to a $\log(n)$ factor. Our results thus uncover a refined trade-off between query smoothness and achievable utility guarantees for differentially private synthetic data.
1.2 Main results
We will consider $(\varepsilon,\delta)$-differential privacy defined below.
Definition 1 ($(\varepsilon,\delta)$-Differential Privacy).
A randomized algorithm $\mathcal{M}:\Omega^{n}\to\mathcal{Z}$ is said to satisfy $(\varepsilon,\delta)$-differential privacy if for all datasets $X,X^{\prime}\in\Omega^{n}$ that differ in at most one entry and for all measurable sets $S\subseteq\mathcal{Z}$,
$\mathbb{P}\!\left(\mathcal{M}(X)\in S\right)\;\leq\;e^{\varepsilon}\,\mathbb{P}\!\left(\mathcal{M}(X^{\prime})\in S\right)\;+\;\delta.$
We focus on the class of smooth queries. For an integer $k\geq 1$, define the $k$-smooth function class
$\mathcal{F}_{k}\coloneqq\Bigl\{\,f:[-1,1]^{d}\to\mathbb{R}\;\Bigm|\;\max_{1\leq|{\boldsymbol{\alpha}}|\leq k}\,\lVert\partial^{{\boldsymbol{\alpha}}}f\rVert_{\infty}\leq 1\,\Bigr\}$ (2)
and the associated integral probability metric [51] between two probability distribution $p,q$ on $1,1]^{d}$:
$d_{k}(p,q)\coloneqq\sup_{f\in\mathcal{F}_{k}}\big|\langle f,p-q\rangle\big|=\sup_{f\in\mathcal{F}_{k}}\left|\int f\mathrm{d}p-\int f\mathrm{d}q\right|.$
When $k=1$, since differentiable functions are dense in the class of 1-Lipschitz functions, this reduces to the $W_{1}$-distance between $p,q$ due to the Kantorovich-Rubinstein duality [57]. For any data set $X=(x_{1},\ldots,x_{n})\in\Omega^{n}$, where $\Omega=[-1,1]^{d}$, we denote its associated empirical distribution as
$p_{X}\coloneqq\frac{1}{n}\sum_{j=1}^{n}\delta_{x_{j}}.$
The utility of a given differentially private synthetic data generation algorithm $\mathcal{M}$ is given by $d_{k}(p_{X},p_{\mathcal{M}(X)})$, where $\mathcal{M}(X)\subseteq\Omega$ is the output of the algorithm $\mathcal{M}$.
Now we are ready to state our algorithmic results:
Theorem 1 (DP synthetic data generation for $k$-smooth queries).
Let $\varepsilon\in(0,1)$. There exists a polynomial-time $(\varepsilon,\delta)$-DP algorithm that outputs a synthetic dataset $Y\subseteq\Omega$ such that
$\mathbb{E}d_{k}(p_{X},p_{Y})\leq\begin{cases}C_{k}d\left(\dfrac{\sqrt{\log(1.25/\delta)}}{\varepsilon n}{}\right)^{k/d}&\text{if }k<d,\\ \dfrac{C_{k}d\sqrt{\log(1.25/\delta)}\log(\varepsilon n)}{\varepsilon n}&\text{if }k=d,\\ \dfrac{C_{k}d\sqrt{\log(1.25/\delta)}}{\varepsilon n}&\text{if }k>d.\end{cases}$ (3)
where $C_{k}\leq(Ck)^{k}$ with some absolute constant $C$.
The detailed DP-synthetic data generation algorithm is presented in Algorithm 1. Theorem 1 uncovers a sharp phase transition in the role of smoothness: once the smoothness order exceeds the dimension ($k>d$), additional smoothness no longer yields further improvements in the error rate. In particular, for multivariate polynomial queries, including low-order moments and correlations, linear statistics, and quadratic functionals, the theorem establishes an error rate of $O_{k}(n^{-1})$, which is independent of the ambient dimension $d$. This strictly improves the $O\left(n^{-\frac{2k}{3d+2k}}\right)$ rate in [60, Theorem 22] when $k=1$, and recovers [48, Theorem 41], since differentiable functions are dense in the class of $1$-Lipschitz functions.
For simplicity, we assume $\varepsilon\in(0,1)$ in Theorem 1. When $\varepsilon>1$, the only modification is the choice of the variance in the Gaussian mechanism step of Algorithm 1. Based on Lemma 1, the same analysis applies, differing only in its dependence on $\varepsilon$ in the utility bound.
The next theorem shows that the rate achieved in 1 is minimax optimal up to a factor $d$ when $k\not=d$ and is optimal up to a logarithmic factor when $k=d$.
Theorem 2 (Minimax lower bound).
For any constant $c_{1}>0$, assume $\varepsilon\in(\frac{1}{n},c_{1}),\delta\in[0,c_{2}(e^{\varepsilon}-1))$ for some constant $c_{2}$ depending only on $c_{1}$. Then the following holds:
$\inf_{\mathcal{M}\text{ is }(\varepsilon,\delta)\text{-DP}}\;\sup_{X\in\Omega^{n}}\mathbb{E}\big[d_{k}(p_{X},p_{\mathcal{M}(X)})\big]\geq C_{k}(n\varepsilon)^{-\min\{1,k/d\}},$ (4)
where the infimum is over all $(\varepsilon,\delta)$-algorithms that output synthetic data $\mathcal{M}(X)\subseteq\Omega$, and $C_{k}$ is a constant depending only on $k$.
The minimax lower bound is stronger and more general compared to the lower bound in [11], where the authors considered only $\varepsilon$-DP algorithms with the utility given by the 1-Wasserstein distance. When taking $\delta=0,k=1$, we recover the $n^{-1/d}$ lower bound in Wasserstein distance from [11]. Our proof is conceptually very different from the packing number argument in [11, Corollary 9.4].
Theorem 2 covers a wide range of $(\varepsilon,\delta)$ used in practice. In particular, $\varepsilon$ can be a constant larger than 1 and $\delta$ can be of order 1. This is in contrast to many existing minimax lower bounds in the literature with restricted ranges of $(\varepsilon,\delta)$. The restriction on $\delta$ in minimax lower bounds is typically not inherent to the definition of $(\varepsilon,\delta)$-DP; rather, it reflects (i) the privacy regime one intends to model and (ii) how the proof composes privacy losses across many individuals, steps, or hypotheses. Broadly, prior techniques fall into two regimes:
  • Vanishing $\delta$ (often $\delta=o(1/n)$). In fingerprinting-code and tracing-based lower bounds for private query answering (e.g., [16, 52]), the argument repeatedly applies approximate DP along a long adaptive process. As a result, the $\delta$-failure probability can accumulate (typically via union bounds) across many steps, and one therefore requires $\delta\ll 1/n$ to keep the overall failure probability negligible. A related requirement appears in the estimation lower-bound framework of [17], where a tracing attack combined with a group-privacy/union-bound step leads to the assumption $\delta=o(1/n)$.
  • $\varepsilon$-relative $\delta$ (no need for $\delta\to 0$). In DP-Le Cam and DP-Assouad [3], approximate DP enters primarily through an additive slack term that scales with the Hamming radius of the reduction (via group privacy), so it suffices to control $\delta$ relative to $\varepsilon$. For example, [3] yields a distinguishability contribution of the form $D\delta$; when the construction takes $D=\Theta(1/\varepsilon)$, assuming $\delta=O(\varepsilon)$ keeps this term bounded. Likewise, [42, Theorem 1.3] assumes $\delta\leq\varepsilon$ so that the accumulated approximate-DP slack remains $O(1)$.
Our lower bound falls into this second regime: we only need privacy leakage to remain uniformly controlled throughout the reduction, and the condition $\delta\lesssim(e^{\varepsilon}-1)$ in Theorem 2 ensures that the accumulated slack does not dominate.
Technical overview
In the work [48], the Chebyshev moment matching method captures information from a dataset $X$ via the Chebyshev moments of $p_{X}$. By perturbing these moments, one can construct a private probability measure $q$. The utility loss $W_{1}(p_{X},q)$ is controlled through the analytic properties of Chebyshev polynomial approximations for 1-Lipschitz functions; in particular, Jackson’s theorem [41] is a key ingredient in achieving the $n^{-1/d}$ rate in [48]. This connection between moment matching and Jackson’s theorem was also recently explored by [5] in the context of empirical measure concentration under the Wasserstein distance.
In the proof of Theorem 1, we leverage the $k$-smoothness of the function class $\mathcal{F}_{k}$ to achieve Chebyshev polynomial approximation rates superior to those provided by Jackson’s theorem for Lipschitz functions. In our setting, a generalized Jackson’s theorem for $k$-smooth functions is needed, and we provide an improved estimate based on [40] tailored to our setting in Lemma 3. In addition, we derive a higher-order global coefficient-decay result for multivariate Chebyshev expansions. Our primary technical contribution is an argument that translates these smoothness assumptions into global energy decay bounds (Theorem 3). While this result is central to our utility analysis of DP-synthetic data in Theorem 1, it may also be of independent interest for broader distribution approximation problems.
Our analysis for $(\varepsilon,\delta)$-DP relies on the $\ell_{2}$-sensitivity analysis based on the Gaussian mechanism, which does not directly apply to the $\ell_{1}$-sensitivity analysis based on the Laplacian mechanism [26]. Achieving an optimal rate for $\varepsilon$-DP synthetic data generation for $k$-smooth queries remains an open question.
In Theorem 2, we establish a minimax lower bound for synthetic data mechanisms by tailoring a DP-Assouad (hypercube) reduction [3] to the $d_{k}$-distance, and the main novelties are the extra ingredients needed to make that reduction both privacy-aware and smoothness-aware. First, we develop an explicit way to turn approximate-DP indistinguishability into a total-variation distance control for outputs when two input datasets differ in many entries. Second, we construct a family of localized smooth bump queries supported on disjoint grid cells, and a matching family of hard datasets. Finally, we connect privacy to utility by a coupling-based Assouad inequality [63]. Optimizing the grid resolution then yields the lower-bound rate and explains the phase transition where additional smoothness stops improving the achievable accuracy.
Organization of the paper
The rest of the paper is organized as follows. We provide preliminary results on differential privacy and Chebyshev polynomials in Section 2. In Section 3, we establish a general theorem connecting the integral probability metric induced by $\mathcal{F}_{k}$ and Chebyshev moments between two probability measures. In Section 4, we present our DP-synthetic data generation algorithm and analyze its privacy and utility guarantees stated in Theorem 1. The proof of the minimax lower bound (Theorem 2) is given in Appendix A. Additional proofs are deferred to the Appendices B to E. The computational complexity of Algorithm 1 is analyzed in Appendix F.
2 Preliminaries
2.1 Differential privacy
The following lemma provides a simple yet useful algorithm to guarantee $(\varepsilon,\delta)$-differential privacy. It suffices to focus on the sensitivity of our target and adding Gaussian noise to the true value.
Lemma 1 (Gaussian Mechanism [26, 7, 58]).
Let $f:\Omega^{n}\to\mathbb{R}^{d}$ with $\ell^{2}$-sensitivity defined as
$\Delta_{2,f}^{2}\coloneqq\sup_{X,X^{\prime}\text{ adjacent}}\|f(X)-f(X^{\prime})\|_{2}^{2}.$
Then the randomized algorithm $\mathcal{M}:X\mapsto f(X)+Z$ with $Z\sim N(0,\sigma^{2}I_{d})$ satisfies $(\varepsilon,\delta)$-differential privacy if
$\sigma=\frac{\sqrt{2}\Delta_{2,f}}{\varepsilon}\cdot\sqrt{\log\frac{1}{4\delta(1-\delta)}+\varepsilon}\quad\text{for }\delta<1/2.$
Moreover, if we assume both $\varepsilon,\delta\in(0,1)$, we can take $\sigma=\frac{\Delta_{2,f}\sqrt{2\log(1.25/\delta)}}{\varepsilon}$.
2.2 Normalized multivariate Chebyshev polynomials
We will use multivariate Chebyshev polynomials in our algorithm and its analysis. Let us start with the 1-dimensional case. On [1,1][-1,1] let
$d\mu(x)=\frac{dx}{\pi\sqrt{1-x^{2}}},$
and on $1,1]^{d}$ the product probability measure $\mu_{d}=\mu^{\otimes d}$. Write ${\boldsymbol{\theta}}=(\theta_{1},\dots,\theta_{d})\in[0,\pi]^{d}$ and $\cos{\boldsymbol{\theta}}=(\cos\theta_{1},\dots,\cos\theta_{d})$. For $n\geq 0$ define the orthonormal Chebyshev basis
$\overline{T}_{0}({\mathbf{x}})\equiv 1;\qquad\overline{T}_{n}({\mathbf{x}})=\sqrt{2}\,\cos(n\arccos x),\ \forall n\geq 1.$
For a multi-index $K=(k_{1},\dots,k_{d})\in\mathbb{N}^{d}$ define the normalized Chebyshev polynomial
$\overline{T}_{K}({\mathbf{x}})=\prod_{i=1}^{d}\overline{T}_{k_{i}}(x_{i}).$
Let $\psi_{0}(\theta)\equiv 1$ and $\psi_{n}(\theta)=\sqrt{2}\cos(n\theta)$ for $n\geq 1$, and set
$\psi_{K}({\boldsymbol{\theta}})=\prod_{i=1}^{d}\psi_{k_{i}}(\theta_{i}).$
All $L^{2}$ norms and inner products of the functions in $\boldsymbol{\theta}$}-space are taken with respect to the probability measure $\pi^{-d}d{\boldsymbol{\theta}}$ on $0,\pi]^{d}$:
$\left\langle u,v\right\rangle:=\frac{1}{\pi^{d}}\int_{[0,\pi]^{d}}u({\boldsymbol{\theta}})\,v({\boldsymbol{\theta}})\,\mathrm{d}{\boldsymbol{\theta}},\quad\left\lVert u\right\rVert_{2}^{2}:=\left\langle u,u\right\rangle,\quad\forall u=u({\boldsymbol{\theta}}),v=v({\boldsymbol{\theta}}).$
The following lemma reveals the relation between the two $L^{2}$ spaces we defined. It indicates that the normalized Chebyshev polynomials $\{\overline{T}_{K}\}$ serve as a role of Fourier basis on [1,1][-1,1] equipped with weight measure $\mu_{d}$. The proof of Lemma 2 is in Appendix B.
Lemma 2 (Isometry to cosine variables and orthonormal basis).
$\{\psi_{K}\}_{K\in\mathbb{N}^{d}}$ is an orthonormal basis of $L^{2}([0,\pi]^{d})$. The linear map $U:L^{2}(\mu_{d})\to L^{2}([0,\pi]^{d},\pi^{-d}\mathrm{d}{\boldsymbol{\theta}})$, $(Uf)({\boldsymbol{\theta}})\coloneqq f(\cos{\boldsymbol{\theta}})$, is an isometry:
$\|Uf\|_{2}^{2}=\|f\|_{L^{2}(\mu_{d})}^{2}.$
Moreover, $U(\overline{T}_{K})=\psi_{K}$ and hence $\{\overline{T}_{K}\}_{K\in\mathbb{N}^{d}}$ is an orthonormal basis of $L^{2}(\mu_{d})$.
3 Integral probability metric and Chebyshev moments
For a multi-index ${\boldsymbol{\alpha}}=(\alpha_{1},\dots,\alpha_{d})$, denote $|{\boldsymbol{\alpha}}|\coloneqq\sum_{i}\alpha_{i}$. For $K=(k_{1},\dots,k_{d})$, denote $\|K\|_{2}^{2}\coloneqq\sum_{i}k_{i}^{2}$. We provide an estimate of the coefficients in the Chebyshev expansion of a $k$-smooth function in $1,1]^{d}$. The proof of Theorem 3 is in Appendix C.
Theorem 3 (Order-$k$ coefficient decay).
Let $d\geq 1$, $k\in\mathbb{N}$, and $f\in C^{k}([-1,1]^{d})\subset L^{2}(\mu_{d})$. With the expansion
$f=\sum_{K\in\mathbb{N}^{d}}c_{K}\,\overline{T}_{K},\qquad c_{K}=\left\langle f,\overline{T}_{K}\right\rangle_{L^{2}(\mu_{d})},$
there is
$\sum_{K}\|K\|_{2}^{2k}\,c_{K}^{2}\;\leq\;C_{k}\,\Big(\max_{1\leq|{\boldsymbol{\beta}}|\leq k}\|\partial_{x}^{\boldsymbol{\beta}}f\|_{L^{\infty}([-1,1]^{d})}\Big)^{2},\text{ where }C_{k}\leq e^{k}k!.$ (5)
We also need the following generalization of Jackson’s Theorem for $k$-smooth functions in $1,1]^{d}$. This can be derived from [40, 19], and we include the proof in Appendix D.
Lemma 3 (Jackson’s Theorem for $d$-dimensional $k$-smooth functions).
Let $d,k\geq 1$ and $f\in\mathcal{F}_{k}$ with expansion $f({\mathbf{x}})=\sum_{K}c_{K}\overline{T}_{K}({\mathbf{x}})$. Then for every integer $m\geq k$, there exists a tensor-product Chebyshev polynomial
$\tilde{f}_{m}({\mathbf{x}})=\sum_{K\in\{0,1,\dots,m-1\}^{d}}\tilde{c}_{K}\overline{T}_{K}({\mathbf{x}}),$
such that
$\|f-\tilde{f}_{m}\|_{\infty}\ \leq\ \frac{dC_{k}^{\mathrm{Jac}}}{m^{k}};\qquad|\tilde{c}_{K}|\leq|c_{K}|,\;\forall K,$
where $C_{k}^{\mathrm{Jac}}\leq(Ck)^{k}$ is a constant depending only on $k$.
Next, we apply the result of the Chebyshev moments of $k$-smooth functions and study the relationship between two types of discrepancies of probability measures $p$ and $q$: the integral probability metric $d_{k}(p,q)$ and the weighted $\ell_{2}$-difference of their Chebyshev moments, stated as follows:
Theorem 4 ($d_{k}$-distance from Chebyshev moments).
For two probability measures $p,q$, let
$\Gamma^{2}\coloneqq\sum_{K\in\{0,...,m\}^{d}\setminus\{0\}}\frac{1}{\|K\|_{2}^{2k}}\left|\mathbb{E}_{X\sim p}\overline{T}_{K}(X)-\mathbb{E}_{X\sim q}\overline{T}_{K}(X)\right|^{2}.$
Then for any integer $k\geq 1$,
$d_{k}(p,q)\coloneqq\sup_{f\in\mathcal{F}_{k}}\big|\langle f,p-q\rangle\big|\leq\frac{2C_{k}^{\mathrm{Jac}}\cdot d}{m^{k}}+\sqrt{C_{k}}\cdot\Gamma,$
where $C_{k}$ is the constant defined in Theorem 3.
Proof.
For any function $\in\mathcal{F}_{k}, ff$ is also $1$-Lipschitz. Then for the function $\tilde{f}=\sum_{K}\tilde{c}_{K}\overline{T}_{K}$, by Jackson’s Theorem introduced in Lemma 3, we have
$\left|\left\langle f,p-q\right\rangle\right|$ $\leq\left|\left\langle\tilde{f},p-q\right\rangle\right|+\left|\left\langle f-\tilde{f},p-q\right\rangle\right|$
$\leq\left|\left\langle\sum_{K}\tilde{c}_{K}\overline{T}_{K}(x),p-q\right\rangle\right|+\left|\left\langle f-\tilde{f},p\right\rangle\right|+\left|\left\langle f-\tilde{f},q\right\rangle\right|$
$\lesssim\sum_{K}|\tilde{c}_{K}|\cdot\left|\left\langle\overline{T}_{K}(x),p-q\right\rangle\right|+\frac{dC_{k}^{\mathrm{Jac}}}{m^{k}}+\frac{dC_{k}^{\mathrm{Jac}}}{m^{k}}.$
For the first term, note that $\left\langle\overline{T}_{K}(x),p-q\right\rangle=\mathbb{E}_{X\sim p}\overline{T}_{K}(X)-\mathbb{E}_{X\sim q}\overline{T}_{K}(X)$ is the difference of the $K$-th Chebyshev moments of $p,q$. By our assumption as well as the Cauchy-Schwarz inequality,
$\sum_{K}|\tilde{c}_{K}|\cdot\left|\left\langle\overline{T}_{K}(x),p-q\right\rangle\right|$ $\leq\sum_{K\in\mathbb{Z}_{\geq 0}^{d}}|c_{K}|\cdot\left|\left\langle\overline{T}_{K}(x),p-q\right\rangle\right|$
$=\sum_{K\neq 0}\|K\|_{2}^{k}|\tilde{c}_{K}|\cdot\frac{\left|\left\langle\overline{T}_{K}(x),p-q\right\rangle\right|}{\|K\|_{2}^{k}}$
$\leq\left(\sum_{K\neq 0}\|K\|_{2}^{2k}|\tilde{c}_{K}|^{2}\right)^{1/2}\left(\sum_{K\neq 0}\frac{\left|\left\langle\overline{T}_{K}(x),p-q\right\rangle\right|^{2}}{\|K\|_{2}^{2k}}\right)^{1/2}\leq\sqrt{C_{k}}\cdot\Gamma,$
where in the last equality we use Theorem 3, which provides the decay rate of the Chebyshev moments of any $k$-th differential function. ∎
4 DP synthetic data generation for $k$-smooth queries
Input: Dataset ${\mathbf{x}}_{1},\dots,{\mathbf{x}}_{n}\in[-1,1]^{d}$, privacy parameters $\varepsilon,\delta>0$, smoothness parameter $k\geq 1$. Moment matching level $m>0$.
1: Let $\mathcal{G}=\{-1,-1+\Delta,-1+2\Delta,\dots,1-\Delta\}^{d}$ be the grid where $\Delta=2\cdot m^{-k}$ is the grid width. Set $r=m^{k}$ and for $J=(j_{1},\dots,j_{d})\in[r]^{d}$ denote
${\mathbf{g}}_{J}=\big(-1+(j_{1}-1)\Delta,\dots,-1+(j_{d}-1)\Delta\big)$
be the $J$-th element of $\mathcal{G}$.
2: For $i=1,\dots,n$, round ${\mathbf{x}}_{i}$ to the nearest point in $\mathcal{G}$:
$\tilde{{\mathbf{x}}}_{i}=\arg\min_{{\mathbf{y}}\in\mathcal{G}}\|{\mathbf{x}}_{i}-{\mathbf{y}}\|_{2}.$
3: Set $\sigma^{2}\geq\frac{4\cdot 2^{d}}{\pi^{d}}\cdot\frac{S\log(1.25/\delta)}{n^{2}\varepsilon^{2}}$, where $S=\sum_{K\in\{0,\dots,m\}^{d}}\frac{1}{\|K\|_{2}^{k}}$.
 For $K\in\{0,\dots,m\}^{d}\setminus\{\mathbf{0}\}$, let $\widehat{m}_{K}=\eta_{K}+\tfrac{1}{n}\sum_{i=1}^{n}\overline{T}_{K}(\tilde{{\mathbf{x}}}_{i})$ with $\eta_{K}\sim N(0,\|K\|_{2}^{k}\sigma^{2})$.
4: Let $\{q_{J}\}_{J\in[r]^{d}}$ be the solution to the following optimization problem:
$\min_{(z_{J})_{J\in[r]^{d}}}\quad$ $\sum_{K\in\{0,\ldots,m\}^{d}\setminus\{\mathbf{0}\}}\frac{1}{\|K\|_{2}^{2k}}\left(\widehat{m}_{K}-\sum_{J\in[r]^{d}}z_{J}\,\overline{T}_{K}({\mathbf{g}}_{J})\right)^{\!2},$
subject to $\sum_{J\in[r]^{d}}z_{J}=1,\qquad z_{J}\geq 0,\ \forall J\in[r]^{d}.$
Output: A synthetic dataset $Y\subseteq[-1,1]^{d}$ according to the probability measure $q=\sum_{J\in[r]^{d}}q_{J}\delta_{g_{J}}$, where $\delta$ is the Dirac delta function.
Algorithm 1 Private Chebyshev moment matching for smooth queries
We now present Algorithm 1, which generalizes the synthetic data algorithm in [48] to $k$-smooth queries. Algorithm 1 is motivated by Theorem 4: by perturbing the Chebyshev moments of the original empirical data distribution, we can obtain a synthetic data distribution with good utility guarantees. The outline of the algorithm can be summarized as follows:
  1. 1.
    Partition the region $1,1]^{d}$ into a grid with resolution $\Delta$.
  2. 2.
    Align the original data to the grid points and get an approximate distribution on the grid.
  3. 3.
    Compute the Chebyshev moments of the approximated distribution and add noise to guarantee privacy.
  4. 4.
    Generated a synthetic data distribution by matching the perturbed Chebyshev moments.
Note that in the output step of Algorithm 1, the size of the synthetic data $|Y|$ depends on the rational approximation of the density function of $q$, and we discuss the details here. Let $q_{J}$ be the weight of the probability measure $q$ on $g_{J}$. We can choose rational numbers $\{w_{J}\}_{J\in[r]^{d}}$ such that $\max_{J}|w_{J}-q_{J}|$ is arbitrarily small. Let $m^{\prime}$ be the least common multiple of the denominators of all $\{w_{J},J\in[r]^{d}\}$, then we output the synthetic dataset $Y$ containing $m^{\prime}w_{J}$ copies of $g_{J}$ for each $J$.
Next, we show the privacy and accuracy guarantee of Algorithm 1, which proves Theorem 1.
4.1 Privacy
For a dataset $X=\{{\mathbf{x}}_{1},\ldots,{\mathbf{x}}_{n}\}\subseteq[-1,1]^{d}$, let $f(X)$ be a vector-valued function with the scaled Chebyshev moments of the uniform distribution over $X$, indexed by $K\in\{0,\ldots,m\}^{d}\setminus\{\mathbf{0}\}$:
$f(X)_{K}=\frac{1}{\|K\|_{2}^{k/2}}\cdot\frac{1}{n}\sum_{i=1}^{n}\overline{T}_{K}({\mathbf{x}}_{i}),$
where $f(X)_{K}$ denotes the $K$-th entry of the vector $f(X)_{K}$, and $\overline{T}_{K}({\mathbf{x}})$ is the $K$-th normalized multivariate Chebyshev polynomial.
Define $\operatorname{nnz}(K)$ to be the number of non-zero entries in $K$. Since $\max_{{\mathbf{x}}\in[-1,1]^{d}}|\overline{T}_{K}({\mathbf{x}})|\leq\sqrt{2^{\mathrm{nnz}(K)}/\pi^{d}}$, for adjacent datasets $X,X^{\prime}$, we have
$|f({X})_{K}-f({X}^{\prime})_{K}|\leq\frac{1}{\|K\|_{2}^{k/2}}\cdot\frac{2}{n}\cdot\max_{{\mathbf{x}}\in[-1,1]^{d}}|\overline{T}_{K}({\mathbf{x}})|\leq\frac{2\sqrt{2^{\mathrm{nnz}(K)}/\pi^{d}}}{\|K\|_{2}^{k/2}\cdot n}.$
Hence, the sensitivity has bound
$\Delta_{2,f}^{2}$ $=\max_{\begin{subarray}{c}{X},{X}^{\prime}\text{ adjacent}\end{subarray}}\|f({X})-f({X}^{\prime})\|_{2}^{2}$
$\leq\sum_{K\in\{0,\ldots,m\}^{d}\setminus\{\mathbf{0}\}}\frac{1}{\|K\|_{2}^{k}}\cdot\frac{1}{n^{2}}\cdot\frac{4\cdot 2^{\mathrm{nnz}(K)}}{\pi^{d}}$
$\leq\frac{4\cdot 2^{d}}{\pi^{d}n^{2}}\sum_{K\in\{0,\ldots,m\}^{d}\setminus\{\mathbf{0}\}}\frac{1}{\|K\|_{2}^{k}}=\frac{4\cdot 2^{d}}{\pi^{d}n^{2}}S,$
where $S=\sum_{K\in\{0,\ldots,m\}^{d}\setminus\{\mathbf{0}\}}\frac{1}{\|K\|_{2}^{k}}$ as defined in Algorithm 1.
Let the $g(X)$ denote the rounded datasets computed Algorithm 1, i.e. $g(X)\coloneqq\{\tilde{\mathbf{x}}_{1},\ldots,\tilde{\mathbf{x}}_{n}\}$. Then for two adjacent datasets ${X},{X}^{\prime}$, we have $g(X),g(X^{\prime})$ are also adjacent. Thus we have
$\Delta_{2,f\circ g}^{2}\leq\Delta_{2,f}^{2}\leq\frac{4\cdot 2^{d}}{\pi^{d}n^{2}}S.$
By Gaussian mechanism (Lemma 1) and our choice of $\sigma^{2}$ in Algorithm 1, adding independent $N(0,\sigma^{2})$ to each component of $f\circ g(x)$ guarantees $(\varepsilon,\delta)$-differential privacy, which is exactly to output $(\|K\|_{2}^{-k/2}\widehat{m}_{K})_{K\in\{0,\ldots,m\}^{d}\setminus\{\mathbf{0}\}}$. Then the moments $(\widehat{m}_{K})_{K\in\{0,\ldots,m\}^{d}\setminus\{\mathbf{0}\}}$ satisfies $(\varepsilon,\delta)$-differential privacy.
Finally, since the remainder of Algorithm 1 simply post-processes $(\widehat{m}_{K})_{K\in\{0,\ldots,m\}^{d}\setminus\{\mathbf{0}\}}$ without returning to the original data $X$, the output of the algorithm is also $(\varepsilon,\delta)$-differentially private.
4.2 Accuracy
Next, we show the accuracy bound of our synthetic data algorithm in the integral probability metric $d_{k}$. Let $\widetilde{{X}}$ denote the aggregated dataset $\{\tilde{x}_{1},\dots,\tilde{x}_{n}\}$ as the output. Then as each data point is moved by at most $\frac{\sqrt{d}}{2}\Delta$, there is
$d_{k}(\mu_{{X}},\mu_{\widetilde{{X}}})\leq\sup_{f:1\text{-Lip}}\big|\langle f,\mu_{{X}}-\mu_{\widetilde{{X}}}\rangle\big|=W_{1}(\mu_{{X}},\mu_{\widetilde{{X}}})\leq\frac{\sqrt{d}}{2}\Delta.$
Here $W_{1}$ denotes the 1-Wasserstein distance, and the equality is from (1). Since $q$ is optimal for construction under the weighted $\ell^{2}$-norm of the Chebyshev moments, there is
$\Gamma^{2}$ $\coloneqq\sum_{K}\frac{\left|\mathbb{E}_{X\sim\mu_{\widetilde{{X}}}}\overline{T}_{K}(X)-\mathbb{E}_{X\sim q}\overline{T}_{K}(X)\right|^{2}}{\|K\|_{2}^{2k}}$
$\leq 2\sum_{K}\frac{\left|\mathbb{E}_{X\sim\mu_{\widetilde{{X}}}}\overline{T}_{K}(X)-\widehat{m}_{K}\right|^{2}+\left|\widehat{m}_{K}-\mathbb{E}_{X\sim q}\overline{T}_{K}(X)\right|^{2}}{\|K\|_{2}^{2k}}$
$\leq 4\sum_{K}\frac{\left|\mathbb{E}_{X\sim\mu_{\widetilde{{X}}}}\overline{T}_{K}(X)-\widehat{m}_{K}\right|^{2}}{\|K\|_{2}^{2k}}=4\sum_{K}\frac{\left|\eta_{K}\right|^{2}}{\|K\|_{2}^{2k}}.$
And by Theorem 4, let $C_{k}$ be the constant in Theorem 3, there is
$d_{k}(\mu_{\widetilde{{X}}},q)\leq\frac{2C_{k}^{\mathrm{Jac}}\cdot d}{m^{k}}+\sqrt{C_{k}}\cdot\Gamma.$
For the term with $\Gamma$, taking expectations,
$(\mathbb{E}\Gamma)^{2}\leq\mathbb{E}\Gamma^{2}=4\sum_{K}\frac{\mathbb{E}\left|\eta_{K}\right|^{2}}{\|K\|_{2}^{2k}}=4\sum_{K}\frac{\sigma^{2}}{\|K\|_{2}^{k}}=\frac{16\cdot 2^{d}}{\pi^{d}}\cdot\frac{\log(1.25/\delta)}{n^{2}\varepsilon^{2}}\cdot S^{2}.$
Applying the estimate of $S$ from Lemma 13 in Appendix E, we have
$\mathbb{E}\Gamma$ $\leq\frac{4\cdot 2^{d/2}}{\pi^{d/2}}\cdot\frac{\sqrt{\log(1.25/\delta)}}{\varepsilon n}\cdot S$ (6)
[…]
where $c>\sqrt{8/\pi}$ is an absolute constant. Therefore, we have an accuracy bound
$\mathbb{E}d_{k}(\mu_{{X}},q)$ $\;\leq\;\mathbb{E}d_{k}(\mu_{{X}},\mu_{\widetilde{{X}}})+\mathbb{E}d_{k}(\mu_{\widetilde{{X}}},q)\;\leq\;\sqrt{d}\cdot\Delta+\frac{2C_{k}^{\mathrm{Jac}}\cdot d}{m^{k}}+\sqrt{C_{k}}\cdot\mathbb{E}\Gamma.$
And in the three different cases, by choosing $m=\Big\lceil\frac{1}{c}\cdot\bigg(\frac{\varepsilon n}{\sqrt{\log(1.25/\delta)}}\bigg)^{1/\max\{d,k\}}\Big\rceil$, we have
  1. 1.
    When $k<d$, $(c\cdot m)^{d}\leq\frac{\varepsilon n}{\sqrt{\log(1.25/\delta)}}$, and there is
    $\mathbb{E}d_{k}(\mu_{{X}},q)$ $\leq\sqrt{d}\cdot\Delta+\frac{2C_{k}^{\mathrm{Jac}}\cdot d}{m^{k}}+\sqrt{C_{k}}\cdot\frac{c^{d}\sqrt{\log(1.25/\delta)}}{\varepsilon n}\cdot m^{d-k}$
    $\leq\;C_{k}^{\prime}d\cdot\bigg(\frac{\sqrt{\log(1.25/\delta)}}{\varepsilon n}\bigg)^{k/d};$
  2. 2.
    When $k=d$, there is
    $\mathbb{E}d_{k}(\mu_{{X}},q)$ $\leq\sqrt{d}\cdot\Delta+\frac{2C_{k}^{\mathrm{Jac}}\cdot d}{m^{k}}+\sqrt{C_{k}}\cdot\frac{c^{d}\sqrt{\log(1.25/\delta)}}{\varepsilon n}\cdot\log m$
    $\leq\;C_{k}^{\prime}d\cdot\frac{\sqrt{\log(1.25/\delta)}}{\varepsilon n}\cdot\log(\varepsilon n);$
  3. 3.
    When $k>d$, there is
    $\mathbb{E}d_{k}(\mu_{{X}},q)$ $\leq\sqrt{d}\cdot\Delta+\frac{2C_{k}^{\mathrm{Jac}}\cdot d}{m^{k}}+\sqrt{C_{k}}\cdot\frac{c^{d}\sqrt{\log(1.25/\delta)}}{\varepsilon n}\cdot 2$
    $\leq\;C_{k}^{\prime}\cdot\frac{\sqrt{\log(1.25/\delta)}}{\varepsilon n}.$
Here we can take $C_{k}^{\prime}=2C_{k}^{\mathrm{Jac}}c^{k}+{2\sqrt{C_{k}}\cdot c^{k}}+1\leq C^{k}{k!}$ for some absolute constant $C>0$.
Acknowledgments
Y.Z. was partially supported by the Simons Grant MPS-TSM-00013944.
Appendix A Minimax lower bound
In this section, we prove Theorem 2 in the following steps.
A.1 Total variance distance bound
We will first introduce several auxiliary lemmas that connect the definition of differential privacy with total variation distance.
Lemma 4 (TV bound from approximate likelihood ratio).
Let $P$ and $Q$ be probability measures on a common measurable space $(\Omega,\mathcal{F})$. Suppose that there exist parameters $\xi\geq 0$ and $\tau\geq 0$ such that for every measurable set $S\in\mathcal{F}$,
$P(S)\ \leq\ e^{\xi}Q(S)+\tau,\qquad Q(S)\ \leq\ e^{\xi}P(S)+\tau.$ (8)
Then
$\operatorname{TV}(P,Q)\coloneqq\sup_{S\in\mathcal{F}}|P(S)-Q(S)|\ \leq\ \tanh\!\Big(\frac{\xi}{2}\Big)+\frac{2\tau}{e^{\xi}+1}.$ (9)
Proof.
We first note that
$\operatorname{TV}(P,Q)=\sup_{S\in\mathcal{F}}|P(S)-Q(S)|=\sup_{S\in\mathcal{F}}\big(P(S)-Q(S)\big),$ (10)
because for any measurable $S$ we have
$P(S^{c})-Q(S^{c})=1-P(S)-\big(1-Q(S)\big)=-\big(P(S)-Q(S)\big).$
Thus, it suffices to bound P(S)Q(S)P(S)-Q(S) from above for any measurable set $S$. For simplicity, fix $S$ and denote
$a\coloneqq P(S),\qquad b\coloneqq Q(S),$
Now applying (8) to $S$ and $S^{c}$ separately, we have that
$P(S)-Q(S)-\tau\ \leq\ (e^{\xi}-1)b,\qquad P(S)-Q(S)-\tau\ \leq\ (e^{\xi}-1)(1-a).$
Combining the above two inequalities,
[…]
Taking the supremum over $S$ finishes the proof. ∎
To control the total variation distance between two output datasets under $(\epsilon,\delta)$-DP, we iterate the DP guarantee along a Hamming path between two datasets as follows:
Lemma 5 (TV distance between two DP-synthetic data outputs).
Let $\mathcal{M}$ be an $(\varepsilon,\delta)$-differentially private mechanism that takes an input dataset in $\Omega^{n}$ and outputs a synthetic dataset. Let $X,Y\in\Omega^{n}$ be two fixed datasets with Hamming distance
$h\ \coloneqq\ d_{\mathrm{Ham}}(X,Y)=\#\{j:x_{j}\neq y_{j}\},$
and denote $Q_{X},Q_{Y}$ as the probability measures of the distributions of the outputs $\mathcal{M}(X),\mathcal{M}(Y)$ correspondingly. Then
$\operatorname{TV}(Q_{X},Q_{Y})\leq\tanh\!\Big(\frac{\varepsilon h}{2}\Big)\Big(1+\frac{2\delta}{e^{\varepsilon}-1}\Big).$ (11)
Proof.
Since $d_{\mathrm{Ham}}(x,y)=h$, there exists a path
$X=X^{(0)},X^{(1)},\dots,X^{(h)}=Y$
in the Hamming graph on $\Omega^{n}$ such that each consecutive pair differs in exactly one coordinate. Write $Q^{(t)}\coloneqq\mathcal{M}\big(X^{(t)}\big)$ for $t=0,\dots,h$. By $(\varepsilon,\delta)$–DP, for every measurable $S$ and every $t$,
$Q^{(t+1)}(S)\ \leq\ e^{\varepsilon}Q^{(t)}(S)+\delta.$
Iterating this inequality for $t=0,1,\dots,h-1$ and unrolling the recursion yields
$Q_{Y}(S)\ =\ Q^{(h)}(S)\ \leq\ e^{\varepsilon h}Q^{(0)}(S)+\delta\sum_{j=0}^{h-1}e^{j\varepsilon}\ =\ e^{\varepsilon h}Q_{X}(S)+\delta_{h},$
where $\delta_{h}=\delta\sum_{j=0}^{h-1}e^{j\varepsilon}\$. Reversing the path from $y$ back to $x$ gives
$Q_{X}(S)\leq e^{\varepsilon h}Q_{Y}(S)+\delta_{h},$ (12)
where
$\delta_{h}=\delta\sum_{j=0}^{h-1}e^{j\varepsilon}=\delta\frac{e^{\varepsilon h}-1}{e^{\varepsilon}-1}.$
Then Lemma 4 yields
$\operatorname{TV}(Q_{X},Q_{Y})\ \leq\ \tanh\!\Big(\frac{\varepsilon h}{2}\Big)+2\delta\left(\frac{e^{\varepsilon h}-1}{e^{\varepsilon}-1}\right)\cdot\frac{1}{e^{\varepsilon h}+1}=\tanh\!\Big(\frac{\varepsilon h}{2}\Big)\Big(1+\frac{2\delta}{e^{\varepsilon}-1}\Big).$
A.2 Construction of a family of test functions
In this section, we construct a collection of test functions in $\mathcal{F}_{k}$ defined in (2), which will later be used in the lower bound proof.
Lemma 6 (A class of test functions in $\mathcal{F}_{k}$).
Fix integers $d\geq 1$ and $k\geq 1$, and let $m\geq 2$. Partition $1,1]^{d}$ into $M=m^{d}$ many uniform closed axis-aligned cubes $\{O_{t}\}_{t=1}^{M}$ of side length $r=\frac{2}{m}$, and let ${\mathbf{a}}_{t}$ be the center of $O_{t}$. For $1\leq t\leq M$, there exists $f_{t}\in C_{c}^{\infty}(O_{t})\cap\mathcal{F}_{k}$ and some point ${\mathbf{x}}_{t}^{\max}\in O_{t}$ such that
$f_{t}({\mathbf{a}}_{t})=0,\quad f_{t}({\mathbf{x}}_{t}^{\max})\geq c_{k}^{\max}r^{k},$
where $c_{k}^{\max}$ is a constant depending only on $k$.
Proof.
Taking a smooth function $\eta\in C_{c}^{\infty}(\mathbb{R})$ such that
$\eta(u)\equiv 1\ \text{for }u\in\Big[-\frac{1}{8},\frac{1}{8}\Big],\qquad\operatorname{supp}(\eta)\subset\Big[-\frac{1}{4},\frac{1}{4}\Big].$
Define a $d$-dimensional bump function
$\chi({\mathbf{u}}):=u_{1}^{k}\prod_{\ell=1}^{d}\eta(u_{\ell}),\qquad{\mathbf{u}}=(u_{1},\dots,u_{d})\in\mathbb{R}^{d},$
and set the normalization constant
$C_{0}:=\max_{1\leq|\alpha|\leq k}\big\|\partial^{\alpha}\chi\big\|_{\infty}\in(0,\infty).$
From our construction of $\chi({\mathbf{u}})$, $C_{0}$ is a constant depending only on $k$. Then, for each $t\in[M]$, define
$f_{t}({\mathbf{x}}):=\frac{r^{k}}{C_{0}}\chi\left(\frac{{\mathbf{x}}-{\mathbf{a}}_{t}}{r}\right).$
(i)Support and disjointness. Since $\operatorname{supp}(\chi)\subset[-\tfrac{1}{4},\tfrac{1}{4}]^{d}$, we have
$\operatorname{supp}(f_{t})\subset{\mathbf{a}}_{t}+r\Big[-\frac{1}{4},\frac{1}{4}\Big]^{d}\subset O_{t}.$
Because these inner boxes lie strictly inside the grid cells, the supports are pairwise disjoint.
(ii) bounds on derivatives For any multi-index $\alpha$ with $1\leq|\alpha|\leq k$, by the chain rule,
$\partial^{\alpha}f_{t}({\mathbf{x}})=\frac{r^{k-|\alpha|}}{C_{0}}\,(\partial^{\alpha}\chi)\!\left(\frac{{\mathbf{x}}-{\mathbf{a}}_{t}}{r}\right).$
Hence
$\|\partial^{\alpha}f_{t}\|_{\infty}\leq\frac{r^{k-|\alpha|}}{C_{0}}\,\|\partial^{\alpha}\chi\|_{\infty}\leq\frac{1}{C_{0}}\max_{|\beta|\leq k}\|\partial^{\beta}\chi\|_{\infty}=1,$
where we used $r\leq 1$ and $k-|\alpha|\geq 0$. Therefore $f_{t}\in\mathcal{F}_{k}$.
(iii) A lower bound at one point. Since $\chi(0)=0$, we have $f_{t}({\mathbf{a}}_{t})=0$. Define
${\mathbf{u}}^{\star}:=\Big(\frac{1}{16},0,\dots,0\Big),\qquad{\mathbf{x}}_{t}^{\max}:={\mathbf{a}}_{t}+r{\mathbf{u}}^{\star}.$
Then, ${\mathbf{u}}^{\star}\in[-\tfrac{1}{8},\tfrac{1}{8}]^{d}$, so $\eta(u^{\star}_{\ell})=1$ for all $\ell$ and thus
$f_{t}({\mathbf{x}}_{t}^{\max})=\frac{r^{k}}{C_{0}}\chi({\mathbf{u}}^{\star})=\frac{r^{k}}{C_{0}}\Big(\frac{1}{16}\Big)^{k}.$
Finally, setting $c^{\max}_{k}=(16^{-k})/C_{0}>0$ gives $f_{t}({\mathbf{x}}_{t}^{\max})\geq c^{\max}_{k}r^{k}$. ∎
A.3 Construction of input datasets
We construct $2^{M}$ many different datasets $X^{\theta},\theta\in\{\pm 1\}^{M}$ as follows:
Construction 1.
Given integers $d\geq 1$, $k\geq 1$ and a grid size $m\geq 2$, let $r\coloneqq 2/m$ and $M\coloneqq m^{d}$. Let $\{O_{t}\}_{t=1}^{M}$ be the uniform partition of $1,1]^{d}$ into closed hypercubes of length $r$ with centers $({\mathbf{a}}_{t})$. Let $\{f_{t}\}_{t=1}^{M}$ be the function class introduced in Lemma 6. Assume $\frac{n}{M}$ is an integer. For any fixed $t\in[M]$, define the multi-sets
$X_{t}^{-1}$ $\coloneqq\{n_{t}\text{ copies of }{\mathbf{a}}_{t}\},$
$X_{t}^{+1}$ $\coloneqq\{n_{t}-\lfloor\beta n\rfloor\text{ copies of }{\mathbf{a}}_{t}\}\cup\{\lfloor\beta n\rfloor\text{ copies of }{\mathbf{x}}_{t}^{\max}\},$
where ${\mathbf{x}}_{t}^{\max}\in O_{t}$ satisfies $f_{t}({\mathbf{x}}_{t}^{\max})\geq c_{d,k}^{\max}r^{k}$, and $\sum_{t=1}^{M}n_{t}=n,\min_{1\leq t\leq M}n_{t}\geq\frac{n}{2M}$. For every binary sequence $\theta\in\{\pm 1\}^{M}$, define a new multi-set $X^{\theta}\coloneqq\bigcup_{t=1}^{M}X_{t}^{\theta_{t}}$ and denote its corresponding empirical measure as $p_{X^{\theta}}$.
The following lemma shows that with the input datasets from Construction 1, if a type of function class exists, this implies a lower bound under the $d_{k}$-distance. Then, analogous to the general DP-Assouad method in [3], we can insert the TV-distance bound into Assouad’s method to control the $d_{k}$-distance.
Lemma 7 (Assouad-type lower bound).
Consider a family of deterministic datasets $\{X^{\theta}\}_{\theta\in\{\pm 1\}^{M}}$ under Construction 1. For a given $\theta\in\{\pm 1\}^{M}$, and $i\in\{1,\dots,M\}$, let $\theta^{(i,+)},\theta^{(i,-)}\in\{\pm 1\}^{M}$ be the two vectors with $\pm 1$ in the $i$-th coordinate, and all other coordinates coincide with $\theta$. Assume that there are functions $\{f_{i}\}_{i=1}^{M}\subseteq\mathcal{F}_{k}$ such that:
  1. 1.
    The supports of $\{f_{i}\}_{i=1}^{M}$ are pairwise disjoint: $\operatorname{supp}(f_{i})\cap\operatorname{supp}(f_{j})=\varnothing$ for $i\neq j$;
  2. 2.
    For all $1\leq i\leq M$, the following integral is independent of $\theta\in\{\pm 1\}^{M}$:
    $\tau_{i}=\int f_{i}({\mathbf{x}})\,\mathrm{d}\big(p_{X^{\theta^{(i,+)}}}-p_{X^{\theta^{(i,-)}}}\big)({\mathbf{x}})\;>\;0.$ (13)
Sample $\theta\sim\mathrm{Unif}(\{\pm 1\}^{M})$. Let $\mathcal{M}$ be any algorithm that maps $X^{\theta}$ to an output dataset $Z\subset\mathcal{X}$. Let $Q_{(i,+)}$ and $Q_{(i,-)}$ denote the conditional distribution of $Z$ conditioning on the events $\{\theta_{i}=+1\}$ and $\{\theta_{i}=-1\}$, respectively. Define
$v_{i}\coloneqq\operatorname{TV}\big(Q_{{(i,+)}},Q_{{(i,+)}}\big).$ (14)
Then
$\mathbb{E}_{\theta,Z}\Big[d_{k}\big(p_{X^{\theta}},p_{Z}\big)\Big]\;\geq\;\frac{1}{2}\sum_{i=1}^{M}(1-v_{i})\tau_{i}.$ (15)
Proof.
For each $i$, define the linear functional
$\Delta_{i}(\theta,Z)\coloneqq\int f_{i}\,\mathrm{d}\big(p_{X^{\theta}}-p_{Z}\big).$
Given any sign vector $s=(s_{1},\dots,s_{M})\in\{\pm 1\}^{M}$, consider $f_{s}\coloneqq\sum_{i=1}^{M}s_{i}f_{i}$. By the disjointness of supports, for every multi-index $\alpha$ with $|\alpha|\leq k$ and every $x\in\mathcal{X}$ there is at most one $i$ such that $x\in\operatorname{supp}(f_{i})$, hence
$\partial^{\alpha}f_{s}(x)=\sum_{i=1}^{M}s_{i}\,\partial^{\alpha}f_{i}(x)=\partial^{\alpha}f_{i_{0}}(x)$
for some $i_{0}$, and therefore $|\partial^{\alpha}f_{s}(x)|\leq\|\partial^{\alpha}f_{i_{0}}\|_{\infty}\leq 1.$ Thus $f_{s}\in\mathcal{F}_{k}$. Then for any $s\in\{\pm 1\}^{M}$, the following lower bound holds:
$d_{k}\big(p_{X^{\theta}},p_{Z}\big)=\sup_{f\in\mathcal{F}_{k}}\left|\int f\,\mathrm{d}\big(p_{X^{\theta}}-p_{Z}\big)\right|\;\geq\;\left|\int f_{s}\,\mathrm{d}\big(p_{X^{\theta}}-p_{Z}\big)\right|=\Big|\sum_{i=1}^{M}s_{i}\,\Delta_{i}(\theta,Z)\Big|.$
For the given pair $(\theta,Z)$, choosing $s_{i}\coloneqq\operatorname{sgn}(\Delta_{i}(\theta,Z))$ (with an arbitrary choice if $\Delta_{i}=0$) gives
$\Big|\sum_{i=1}^{M}s_{i}\,\Delta_{i}(\theta,Z)\Big|=\sum_{i=1}^{M}|\Delta_{i}(\theta,Z)|,$
and we obtain
$d_{k}\big(p_{X^{\theta}},p_{Z}\big)\;\geq\;\sum_{i=1}^{M}|\Delta_{i}(\theta,Z)|.$ (16)
Next, we will bound the right-hand side of (16) in expectation. Fix index $i$. For simplicity, denote
$\int f_{i}\,\mathrm{d}p_{X^{\theta^{(i,+)}}}\eqqcolon a,\qquad\int f_{i}\,\mathrm{d}p_{X^{\theta^{(i,-)}}}\eqqcolon b,\qquad a-b=\tau_{i}>0,$
and
$T_{i}(Z)\coloneqq\int f_{i}\,\mathrm{d}p_{Z}.$
Recall the following conditional distribution we defined:
$Z\mid(\theta_{i}=+1)\sim Q_{{(i,+)}},\qquad Z\mid(\theta_{i}=-1)\sim Q_{{(i,-)}},$
and $v_{i}=\operatorname{TV}\big(Q_{{(i,+)}},Q_{{(i,+)}}\big)$. Let $(Z_{(i,+)},Z_{(i,-)})$ be a maximal coupling of $Q_{{(i,+)}}$ and $Q_{{(i,-)}}$ so that
$\mathbb{P}(Z_{(i,+)}=Z_{(i,-)})=1-v_{i}.$
On the event $\{Z_{(i,+)}=Z_{(i,-)}\}$, we have $T_{i}(Z_{(i,+)})=T_{i}(Z_{(i,-)})$ hence
$\frac{1}{2}\left(\Big|a-T_{i}(Z_{(i,+)})\Big|+\Big|b-T_{i}(Z_{(i,-)})\Big|\right)\;\geq\;\frac{|a-b|}{2}\,\mathbf{1}_{\{Z_{(i,+)}=Z_{(i,-)}\}}\;=\;\frac{\tau_{i}}{2}\,\mathbf{1}_{\{Z_{(i,+)}=Z_{(i,-)}\}}.$
Taking expectations with respect to the coupling gives
$\frac{1}{2}\,\mathbb{E}\left[\Big|a-T_{i}(Z_{(i,+)})\Big|+\Big|b-T_{i}(Z_{(i,-)})\Big|\right]\;\geq\;\frac{\tau_{i}}{2}(1-v_{i}).$
Then we can take expectations in (16). Since $\theta_{i}\sim\mathrm{Uniform}\{\pm 1\}$, by revealing the $i$-th coordinate of $\theta$ first, we further obtain
$\mathbb{E}_{\theta,Z}|\Delta_{i}(\theta,Z)|$ $=\frac{1}{2}\mathbb{E}_{Z\sim Q^{(i,+)}}\Big|a-T_{i}(Z)\Big|+\frac{1}{2}\mathbb{E}_{Z\sim Q^{(i,-)}}\Big|b-T_{i}(Z)\Big|$ (17)
$=\frac{1}{2}\mathbb{E}\left(\Big|a-T_{i}(Z_{(i,+)})\Big|+\Big|b-T_{i}(Z_{(i,-)})\Big|\right)\;\geq\;\frac{\tau_{i}}{2}(1-v_{i}).$ (18)
Summing over $i$, we obtain
$\mathbb{E}_{\theta,Z}\Big[d_{k}\big(p_{X^{\theta}},p_{Z}\big)\Big]\;\geq\;\sum_{i=1}^{M}\mathbb{E}_{\theta,Z}\big|\Delta_{i}(\theta,Z)\big|\;\geq\;\frac{1}{2}\sum_{i=1}^{M}(1-v_{i})\,\tau_{i},$
as desired. ∎
A.4 Proof of Theorem 2
Putting all ingredients together, we will finish the proof of Theorem 2 based on Lemma 6, Lemma 7, and datasets from Construction 1.
Proof of Theorem 2.
Let $\mathcal{M}$ be any $(\varepsilon,\delta)$–DP algorithm that outputs synthetic data in $1,1]^{d}$. Let $\theta\sim\mathrm{Unif}(\{\pm 1\}^{M})$, $M=m^{d}$, $Z\coloneqq\mathcal{M}(X^{\theta})$, and $p_{Z}$ be its associated empirical distribution. Let $Q_{(i,+)}$ and $Q_{(i,-)}$ be the distribution of $Z$ conditioning on the events $\{\theta_{i}=+1\}$ and $\{\theta_{i}=-1\}$, correspondingly, and let
$v_{i}\coloneqq\operatorname{TV}\big(Q_{{(i,+)}},Q_{{(i,-)}}\big),\quad 1\leq i\leq M.$
Step 1: Uniform upper bound for $v_{i}$. For any $\theta\in\{-1,1\}^{M}$, we associate a pair of vectors $(\theta^{(i,+)},\theta^{(i,-)})$ that differs only on the $i$-th coordinate and coincide with $\theta$ for all other coordinates.
For any constant $c_{1}>0$ and $\varepsilon\in(0,c_{1})$, choosing $\beta=\frac{2c_{1}}{n\varepsilon}$ in Construction 1, we have
$d_{\mathrm{Ham}}(X^{\theta^{(i,+)}},X^{\theta^{(i,-)}})\leq\lfloor\beta n\rfloor\leq 2c_{1}/\varepsilon.$
Let $Q_{X^{\theta^{(i,+)}}},Q_{X^{\theta^{(i,-)}}}$ be the distribution of $\mathcal{M}(X^{\theta^{(i,+)}}),\mathcal{M}(X^{\theta^{(i,-)}})$. Then by Lemma 5, under the assumption
$0\leq\varepsilon\leq c_{1},\quad\delta\leq c_{2}(e^{\varepsilon}-1),\quad\text{where}\quad 0<c_{2}<\frac{1}{2}\left(\frac{1}{\tanh(c_{1})}-1\right),$
we have
$\operatorname{TV}(Q_{X^{\theta^{(i,+)}}},Q_{X^{\theta^{(i,-)}}})$ $\leq\tanh\left(\frac{\varepsilon\lfloor\beta n\rfloor}{2}\right)(1+\frac{2\delta}{e^{\varepsilon}-1})$ (19)
$\leq\tanh(c_{1})(1+\frac{2\delta}{e^{\varepsilon}-1})$ (20)
$<\tanh(c_{1})(1+2c_{2})<1.$ (21)
Denote the constant $c_{3}:=\tanh(c_{1})(1+2c_{2})<1$. Since $Q_{(i,+)},Q_{(i,-)}$ is the weighted sum of all probability measures $Q_{X^{\theta^{(i,+)}}},Q_{X^{\theta^{(i,-)}}}$ over all but $i$-th coordinate,
$v_{i}=\operatorname{TV}\big(Q_{{(i,+)}},Q_{{(i,-)}}\big)\leq\frac{1}{2^{M-1}}\!\!\sum_{(\theta^{(i,+)},\theta^{(i,-)})}\operatorname{TV}(Q_{X^{\theta^{(i,+)}}},Q_{X^{\theta^{(i,-)}}})\leq c_{3}<1,$
where the sum is over all pairs of $\{(\theta^{(i,+)},\theta^{(i,-)})\},\theta\in\{\pm 1\}^{M}$ that differs only on the $i$-th coordinate.
Step 2: Apply Lemma 7. We next check our function class $\{f_{t}\}_{t=1}^{M}$ introduced in Lemma 6 satisfies the three conditions from Lemma 7.
Conditions (i), (ii) hold by Lemma 6. For Condition (iii), note that for any pair of $(\theta^{(i,+)},\theta^{(i,-)})$ that differs only on the $i$-th coordinate, by Construction 1, the differences between datasets $X^{\theta^{(i,+)}}$ and $X^{\theta^{(i,-)}}$ are exactly $\lfloor\beta n\rfloor$ many points at either ${\mathbf{x}}_{i}^{\max}$ or ${\mathbf{a}}_{i}$, which is independent of the values of $\theta^{(i,+)},\theta^{(i,-)}$ on the other coordinates. Hence,
$\tau_{i}=\int f_{i}\,\mathrm{d}\big(p_{X^{\theta^{(i,+)}}}-p_{X^{\theta^{(i,-)}}}\big)=\frac{\lfloor\beta n\rfloor}{n}\left(f_{i}({\mathbf{x}}_{i}^{\max})-f_{i}({\mathbf{a}}_{i})\right)$
is a constant independent of $\theta$. Since $\beta=\frac{2c_{1}}{n\varepsilon}$, and $0<\varepsilon\leq c_{1}$, from Lemma 6,
$\tau_{i}\geq\frac{\beta n-1}{n}\left(c_{k}^{\max}r^{k}-0\right)\geq\frac{\beta}{2}c_{k}^{\max}r^{k}.$ (22)
Recall $r=2/m$ from Construction 1. This implies
$\tau_{i}\geq C_{k}\beta m^{-k}$
for some constant $C_{k}>0$ depending on $k$. Plugging the bounds for $v_{i},\tau_{i}$ into (15) in Lemma 7, and using $M=m^{d}$, we obtain
$\mathbb{E}_{\theta,Z}\big[d_{k}(p_{\theta},p_{Z})\big]\;\geq\;MC_{k}\beta m^{-k}\cdot(1-c_{3})\gtrsim C_{k}\frac{m^{d-k}}{n\varepsilon}.$ (23)
Step 3: Optimize over the grid size $m$. Since there are $n_{t}\geq\frac{n}{2M}$ many data points in each cell $O_{t}$ from Construction 1, the requirement that each cell contains at least $\lfloor\beta n\rfloor$ movable points is $\beta n\leq n/m^{d}$. Since $\beta=\frac{2c_{1}}{n\varepsilon}$, this reads $m^{d}\lesssim n\varepsilon$. Thus all admissible $m$ satisfy $1\leq m\lesssim(n\varepsilon)^{1/d}$. Define
$\Phi(m)\coloneqq m^{d-k}/(n\varepsilon).$
If $d>k$ then $\Phi(m)$ is increasing in $m$, so the optimum is attained at $m\asymp(n\varepsilon)^{1/d}$ and
$\Phi(m)\asymp\frac{(n\varepsilon)^{(d-k)/d}}{n\varepsilon}=(n\varepsilon)^{-k/d}.$
If $d=k$, then $\Phi(m)\equiv(n\varepsilon)^{-1}$ is independent of $m$. If $d<k$, then $\Phi(m)$ is decreasing in $m$, so the optimum is attained at the smallest admissible $m$ (a fixed constant), which again yields $\Phi(m)\asymp(n\varepsilon)^{-1}$. Optimizing in $m$ for (23) yields
$\mathbb{E}_{\theta,Z}\big[d_{k}(p_{\theta},p_{Z})\big]\;\gtrsim\;\begin{cases}C_{k}(n\varepsilon)^{-k/d},&d>k,\\ C_{k}(n\varepsilon)^{-1},&d\leq k.\end{cases}$ (24)
Finally, note that
$\sup_{X\in\Omega^{n}}\mathbb{E}\big[d_{k}(\mathrm{Unif}(X),\mathcal{M}(X))\big]\;\geq\;\mathbb{E}_{\theta}\mathbb{E}\big[d_{k}(p_{X^{\theta}},\mathcal{M}(X^{\theta}))\big]=\mathbb{E}_{\theta,Z}\big[d_{k}(p_{\theta},p_{Z})\big],$
so the lower bound (24) for the Bayes risk with respect to the distribution of $\theta$ also lower bounds the minimax risk. Taking an infimum over all $(\varepsilon,\delta)$-differentially private algorithm $\mathcal{M}$ finishes the proof. ∎
Appendix B Proof of Lemma 2
Proof.
Since $dx_{i}=-\sin\theta_{i}\,d\theta_{i}$ and $\sqrt{1-x_{i}^{2}}=\sin\theta_{i}$ for $\theta_{i}\in[0,\pi]$,
$\int_{[-1,1]^{d}}|f({\mathbf{x}})|^{2}\,d\mu_{d}({\mathbf{x}})=\frac{1}{\pi^{d}}\int_{[0,\pi]^{d}}|f(\cos{\boldsymbol{\theta}})|^{2}\,d{\boldsymbol{\theta}}=\|Uf\|_{2}^{2}.$
For $n\geq 1$, $U(\overline{T}_{n})({\boldsymbol{\theta}})=\sqrt{2}\cos(n{\boldsymbol{\theta}})=\psi_{n}({\boldsymbol{\theta}})$ and $U(\overline{T}_{0})=1=\psi_{0}$. The orthogonality and completeness follow from the computation of the one-dimensional case. ∎
Appendix C Proof of Theorem 3
For a multi-index ${\boldsymbol{\alpha}}=(\alpha_{1},\dots,\alpha_{d})$, denote
$|{\boldsymbol{\alpha}}|\coloneqq\sum_{i}\alpha_{i},\quad{\boldsymbol{\alpha}}!\coloneqq\prod_{i}\alpha_{i}!$
and for $K=(k_{1},\dots,k_{d})$, denote $K^{2{\boldsymbol{\alpha}}}\coloneqq\prod_{i=1}^{d}k_{i}^{2\alpha_{i}}$, $\|K\|_{2}^{2}\coloneqq\sum_{i}k_{i}^{2}$.
We prove Theorem 3 with several lemmas. The outline of the proof of Theorem 3 is to show the following inequalities:
$\sum_{K}\|K\|_{2}^{2k}\,c_{K}^{2}\;\leq\;\sum_{K}K^{2{\boldsymbol{\alpha}}}c_{K}^{2}\;=\;\|\partial_{\boldsymbol{\theta}}^{\boldsymbol{\alpha}}g\|_{2}^{2}\;\leq\;C_{d,k}\,\Big(\max_{|{\boldsymbol{\beta}}|\leq k}\|\partial_{x}^{\boldsymbol{\beta}}f\|_{L^{\infty}([-1,1]^{d})}\Big)^{2}.$
As introduced in Lemma 2, the Chebyshev polynomials corresponds to Fourier basis after a step of change-of-variable. Therefore we will first study the $k$-smooth functions in $L^{2}([0,\pi],\pi^{-d}\mathrm{d}{\boldsymbol{\theta}})$. The following lemma is a classical result in Fourier analysis, proven by integration by parts. (See [54, Section 3, Eq (1.4), (1.23)] for more details.)
Lemma 8 (Parseval with derivatives in $\boldsymbol{\theta}$}-space, [54]).
With $g({\boldsymbol{\theta}})=f(\cos{\boldsymbol{\theta}})$ and $g=\sum_{K}c_{K}\psi_{K}\in L^{2}([0,\pi]^{d})$, for each ${\boldsymbol{\alpha}}\in\mathbb{N}^{d}$, $|{\boldsymbol{\alpha}}|\leq k$, $\partial^{{\boldsymbol{\alpha}}}_{g}g$ is continuous and
$\|\partial_{\boldsymbol{\theta}}^{\boldsymbol{\alpha}}g\|_{2}^{2}=\sum_{K}K^{2{\boldsymbol{\alpha}}}c_{K}^{2}.$
Lemma 9 (Chain rule with bounded trigonometric coefficients).
Fix $k\in\mathbb{N}$. For every ${\boldsymbol{\alpha}}\in\mathbb{N}^{d}$ with $|{\boldsymbol{\alpha}}|\leq k$, there exist trigonometric polynomials $P_{{\boldsymbol{\alpha}},{\boldsymbol{\beta}}}({\boldsymbol{\theta}})$, indexed by multi-indices $\boldsymbol{\beta}$} with $0\leq{\boldsymbol{\beta}}\leq{\boldsymbol{\alpha}}$ (coordinate-wise), such that
$\partial_{\boldsymbol{\theta}}^{\boldsymbol{\alpha}}g({\boldsymbol{\theta}})=\sum_{0\leq{\boldsymbol{\beta}}\leq{\boldsymbol{\alpha}}}P_{{\boldsymbol{\alpha}},{\boldsymbol{\beta}}}({\boldsymbol{\theta}})\,\partial_{\mathbf{x}}^{\boldsymbol{\beta}}f(\cos{\boldsymbol{\theta}}),$ (25)
each $P_{{\boldsymbol{\alpha}},{\boldsymbol{\beta}}}$ has degree at most $|{\boldsymbol{\alpha}}|$ in every $\theta_{i}$, and
$\sup_{{\boldsymbol{\theta}}\in[0,\pi]^{d}}|P_{{\boldsymbol{\alpha}},{\boldsymbol{\beta}}}({\boldsymbol{\theta}})|\ \leq|{\boldsymbol{\alpha}}|!\qquad\forall\;0\leq{\boldsymbol{\beta}}\leq{\boldsymbol{\alpha}}.$ (26)
To show the infinity norm bound in Lemma 9, we introduce the Bernstein inequality in harmonic analysis, which states that the infinity norm of a triangular polynomial can be controlled by its derivative.
Lemma 10 (A Bernstein inequality for trigonometric polynomials, [43], Ex. 7.15(c), (7.25)).
Let $Q(\vartheta)=\sum_{n=-m}^{m}a_{n}e^{in\vartheta}$ be a trigonometric polynomial of degree $\leq m$. Then
$\sup_{\vartheta\in\mathbb{R}}|Q^{\prime}(\vartheta)|\ \leq\ m\sup_{\vartheta\in\mathbb{R}}|Q(\vartheta)|.$
Proof of Lemma 9.
Induct on $|{\boldsymbol{\alpha}}|$. For ${\boldsymbol{\alpha}}=0$, we have $P_{0,0}\equiv 1$.
Suppose (25) holds for some $\boldsymbol{\alpha}$}. Next we fix $j\in\{1,\dots,d\}$ and consider ${\boldsymbol{\alpha}}+{\mathbf{e}}_{j}$ where ${\mathbf{e}}_{j}$ is the vector with one at $j$-th entry and zero elsewhere. Applying $\partial_{\theta_{j}}$ we have
$\partial_{\theta_{j}}\partial_{\boldsymbol{\theta}}^{\boldsymbol{\alpha}}g=\sum_{0\leq{\boldsymbol{\beta}}\leq{\boldsymbol{\alpha}}}\big(\partial_{\theta_{j}}P_{{\boldsymbol{\alpha}},{\boldsymbol{\beta}}}\big)\,\partial_{x}^{\boldsymbol{\beta}}f(\cos{\boldsymbol{\theta}})-\sum_{0\leq{\boldsymbol{\beta}}\leq{\boldsymbol{\alpha}}}(\sin\theta_{j})P_{{\boldsymbol{\alpha}},{\boldsymbol{\beta}}}({\boldsymbol{\theta}})\,\partial_{x}^{{\boldsymbol{\beta}}+{\mathbf{e}}_{j}}f(\cos{\boldsymbol{\theta}}),$
since $\partial_{\theta_{j}}=\!-\sin\theta_{j}\,\partial_{x_{j}}$. Relabel ${\boldsymbol{\beta}}\mapsto{\boldsymbol{\beta}}-{\mathbf{e}}_{j}$ in the second sum (interpreting $P_{{\boldsymbol{\alpha}},{\boldsymbol{\beta}}-{\mathbf{e}}_{j}}\equiv 0$ if $\beta_{j}=0$) to obtain the representation for ${\boldsymbol{\alpha}}+{\mathbf{e}}_{j}$ with
$P_{{\boldsymbol{\alpha}}+e_{j},{\boldsymbol{\beta}}}=\partial_{\theta_{j}}P_{{\boldsymbol{\alpha}},{\boldsymbol{\beta}}}-(\sin\theta_{j})P_{{\boldsymbol{\alpha}},{\boldsymbol{\beta}}-{\mathbf{e}}_{j}}\qquad\forall\;0\leq{\boldsymbol{\beta}}\leq{\boldsymbol{\alpha}}.$
From the recursion formula, $P_{{\boldsymbol{\alpha}},{\boldsymbol{\beta}}}$’s are trigonometric polynomials of degree $\leq|{\boldsymbol{\alpha}}|$ in each coordinate $\theta_{j}$.
For the uniform upper bound, define $M_{t}\coloneqq\max\{\sup_{\boldsymbol{\theta}}|P_{{\boldsymbol{\alpha}},{\boldsymbol{\beta}}}({\boldsymbol{\theta}})|:|{\boldsymbol{\alpha}}|=t,\,0\leq{\boldsymbol{\beta}}\leq{\boldsymbol{\alpha}}\}.$ Using Lemma 10 in the $j$-th variable and $|\sin\theta_{j}|\leq 1$,
$\sup_{\boldsymbol{\theta}}|P_{{\boldsymbol{\alpha}}+{\mathbf{e}}_{j},{\boldsymbol{\beta}}}|\leq\sup_{\boldsymbol{\theta}}|\partial_{\theta_{j}}P_{{\boldsymbol{\alpha}},{\boldsymbol{\beta}}}|+\sup_{\boldsymbol{\theta}}|P_{{\boldsymbol{\alpha}},{\boldsymbol{\beta}}-e_{j}}|\leq|{\boldsymbol{\alpha}}|M_{|{\boldsymbol{\alpha}}|}+M_{|{\boldsymbol{\alpha}}|}\leq(|{\boldsymbol{\alpha}}|{+}1)M_{|{\boldsymbol{\alpha}}|}.$
This yields $M_{t+1}\leq(t+1)\,M_{t}$, and $M_{0}=1$, hence $M_{t}\leq t!$. This proves (26).
Proposition 1 (Uniform bound for $\|\partial_{\boldsymbol{\theta}}^{\boldsymbol{\alpha}}g\|_{L^{2}([0,\pi]^{d})}$).
For any $f\in C^{k}([-1,1]^{d})$, $g=f(\cos{\boldsymbol{\theta}})$, and index $\boldsymbol{\alpha}$} with $|{\boldsymbol{\alpha}}|\leq k$,
$\|\partial_{\boldsymbol{\theta}}^{\boldsymbol{\alpha}}g\|_{L^{2}([0,\pi]^{d})}\leq|{\boldsymbol{\alpha}}|!\,\left(1+\frac{k}{d}\right)^{d}\,\max_{|{\boldsymbol{\beta}}|\leq k}\|\partial_{x}^{\boldsymbol{\beta}}f\|_{L^{\infty}([-1,1]^{d})}.$
Proof.
Using the expansion formula (25), we have
$\|\partial_{\boldsymbol{\theta}}^{\boldsymbol{\alpha}}g\|_{L^{2}([0,\pi]^{d})}$ $\leq\|\partial_{\boldsymbol{\theta}}^{\boldsymbol{\alpha}}g\|_{L^{\infty}([0,\pi]^{d})}$
$\leq\sum_{0\leq{\boldsymbol{\beta}}\leq{\boldsymbol{\alpha}}}\sup_{\boldsymbol{\theta}}|P_{{\boldsymbol{\alpha}},{\boldsymbol{\beta}}}({\boldsymbol{\theta}})|\cdot\|\partial_{x}^{\boldsymbol{\beta}}f(\cos{\boldsymbol{\theta}})\|_{L^{\infty}([0,\pi]^{d})}$
$\leq\sum_{0\leq{\boldsymbol{\beta}}\leq{\boldsymbol{\alpha}}}\sup_{\boldsymbol{\theta}}|P_{{\boldsymbol{\alpha}},{\boldsymbol{\beta}}}({\boldsymbol{\theta}})|\cdot\|\partial_{x}^{\boldsymbol{\beta}}f\|_{L^{\infty}([-1,1]^{d})}.$
Note that here the last inequality is due to our $L^{2}$ measure being normalized. As the number of $\boldsymbol{\beta}$} with $0\leq{\boldsymbol{\beta}}\leq{\boldsymbol{\alpha}}$ is $\prod_{i=1}^{d}(\alpha_{i}+1)\leq(1+\frac{k}{d})^{d}$ by Jensen’s inequality. Applying the uniform bound (26) concludes the proof. ∎
Proof of Theorem 3.
We will show
$\sum_{K}\|K\|_{2}^{2k}\,c_{K}^{2}\;\leq\;\sum_{K}K^{2{\boldsymbol{\alpha}}}c_{K}^{2}\;=\;\|\partial_{\boldsymbol{\theta}}^{\boldsymbol{\alpha}}g\|_{2}^{2}\;\leq\;C_{d,k}\,\Big(\max_{|{\boldsymbol{\beta}}|\leq k}\|\partial_{x}^{\boldsymbol{\beta}}f\|_{L^{\infty}([-1,1]^{d})}\Big)^{2}$ (27)
to prove the theorem. Note that the second step in (27) is showed in Lemma 8 and the last step is proved in Proposition 1 with
$C_{k}=k!\left(1+\frac{k}{d}\right)^{d}\leq k!\cdot e^{k}.$
It remains to show the first inequality in (27). For any given $K$ and $a_{i}\coloneqq k_{i}^{2}\geq 0$, we expand
$\Big(\sum_{i=1}^{d}a_{i}\Big)^{k}=\sum_{|{\boldsymbol{\alpha}}|=k}\frac{k!}{{\boldsymbol{\alpha}}!}a^{\boldsymbol{\alpha}}\leq k!\sum_{|{\boldsymbol{\alpha}}|=k}a^{\boldsymbol{\alpha}}\quad.$
Multiplying by a factor $c_{K}^{2}$ and summing over $K$ gives the first inequality in (27). ∎
Appendix D Multivariate Jackson’s Theorem for $k$-smooth functions
To show the multivariate Jackson’s Theorem, we first consider the measure $\pi^{-1}\mathrm{d}x$ on $0,\pi]$ and introduce a powerful tool in the construction:
Lemma 11 (Jackson kernel, [40]).
For integer $m,k\geq 1$, there exists a even kernel function $J_{1}(x):\mathbb{R}\to\mathbb{R}$ with the following properties:
  1. 1.
    (Triangular polynomial) $J_{1}(x)=\sum_{i=0}^{m}\widehat{J}_{1}(i)\cos(ix)$;
  2. 2.
    (Positivity) $J_{1}(x)\geq 0$;
  3. 3.
    (Normalization) $\frac{1}{\pi}\int_{0}^{\pi}J_{1}(x)\mathrm{d}x=1$;
  4. 4.
    (Bounded moments) $\frac{1}{\pi}\int_{0}^{\pi}|x|^{\ell}J_{1}(x)\mathrm{d}x\leq\frac{C}{m^{\ell}}$ for $\ell\leq k$.
Lemma 12 (Jackson’s Theorem for 1-dimensional $k$-smooth functions, [40]).
Let $m,k\geq 1$ and a even periodic function $g$ with period $2\pi$ satisfies $\|g^{(k)}\|_{L^{\infty}([0,\pi]^{d})}\leq 1$. Then for every integer $m\geq k$,
$\tilde{g}_{m}(x)\coloneqq g*J_{1}=\frac{1}{\pi}\int_{0}^{\pi}J_{1}(y)g(x-y)\mathrm{d}y,$
satisfies
$\|g-\tilde{g}_{m}\|_{L^{\infty}([0,\pi]^{d})}\ \leq\ \frac{C_{1}^{k}}{m^{k}}.$
Moreover, if we write the Fourier expansion
$g(x)=\sum_{i=0}^{\infty}c_{i}\cos(ix),\qquad\tilde{g}_{m}(x)=\sum_{i=0}^{m}\tilde{c}_{i}\cos(ix)$
then $|\tilde{c}_{i}|\leq|c_{i}|,\;\forall\;0\leq i\leq k.$
Proof.
The approximation in $\|\cdot\|_{L^{\infty}([0,\pi]^{d})}$ norm is already proved in [40] with the properties in Lemma 11. Here, we simply complete the proof for the bound on the coefficients. For the Jackson’s kernel $J_{1}$, using the property in Lemma 11, we have for $i\geq 0$
$|\tilde{c}_{i}|=|c_{i}|\cdot\left|\frac{1}{\pi}\int_{0}^{\pi}J_{1}(x)\cdot\cos(ix)\mathrm{d}x\right|\leq|c_{i}|\cdot\frac{1}{\pi}\int_{0}^{\pi}J_{1}(x)\mathrm{d}x=|c_{i}|.$
We are now ready to extend the result to general $d$-dimensional space and prove Lemma 3.
Proof of Lemma 3.
For any even periodic function $G({\boldsymbol{\theta}})=G(\theta_{1},\dots,\theta_{d})$, define operator $\mathcal{A}_{i}$ such that $\mathcal{A}_{i}(G)$ is to apply Lemma 12 on the variable $\theta_{i}$:
$\mathcal{A}_{i}(G)(\theta_{1},\dots,\theta_{d})=\frac{1}{\pi}\int_{0}^{\pi}J_{1}(y)G(\theta_{1},\dots,\theta_{i-1},\theta_{i}-y,\theta_{i+1},\dots\theta_{d})\mathrm{d}y.$
Then $\mathcal{A}_{i}(G)$ is a trigonometric polynomial of degree $m$ in $\theta_{i}$. Note that $\mathcal{A}_{i}$ is linear. Therefore, if $g$ is a trigonometric polynomial of degree $m$ in $\theta_{j}$, then $\mathcal{A}_{i}(g)$ is a trigonometric polynomial of degree $m$ in both $\theta_{i},\theta_{j}$. Moreover, since $J_{1}$ is normalized, we have
$\|\mathcal{A}_{i}(G)\|_{L^{\infty}([0,\pi]^{d})}\leq\|G\|_{L^{\infty}([0,\pi]^{d})}\cdot\frac{1}{\pi}\int_{0}^{\pi}J_{1}(x)\mathrm{d}x=\|G\|_{\infty}.$
First, we will consider $g({\boldsymbol{\theta}})=f(\cos{\boldsymbol{\theta}})$ as an even periodic function in each coordinate $\theta_{i}$. We will run the approximation process in each coordinate. Define $\tilde{g}({\boldsymbol{\theta}})=\mathcal{A}_{d}(\mathcal{A}_{d-1}(\cdots\mathcal{(}A_{1}(g({\boldsymbol{\theta}}))))$. By our observation above, $\tilde{g}$ is a trigonometric polynomial of degree $m$ in all $\theta_{i}$, $1\leq i\leq d$. Hence, we can write
$\tilde{g}=\sum_{K\in\{0,\dots,m\}^{d}}\tilde{c}_{K}\prod_{i=1}^{d}\cos(k_{i}\theta_{i}).$
And we aim to show the $\|\cdot\|_{L^{\infty}([0,\pi]^{d})}$ approximation and the entry-wise bound for $\tilde{c}_{K}$.
By Lemma 12, for any even periodic function $G({\boldsymbol{\theta}})$, there is
$\|\mathcal{A}_{i}(G)-G\|_{L^{\infty}([0,\pi]^{d})}\leq\frac{C_{1}^{k}}{m^{k}}\cdot\|\partial_{i}^{k}G\|_{L^{\infty}([0,\pi]^{d})}$
and
$\|\partial_{j}^{k}\mathcal{A}_{i}(G)\|_{L^{\infty}([0,\pi]^{d})}=\|\mathcal{A}_{i}(\partial_{j}^{k}G)\|_{L^{\infty}([0,\pi]^{d})}\leq\|\partial_{j}^{k}G\|_{L^{\infty}([0,\pi]^{d})},\quad\text{when }i\neq j.$
Therefore, by a telescoping sum argument, we will get
$\|Q-g\|_{L^{\infty}([0,\pi]^{d})}\leq\sum_{i=1}^{d}\frac{C_{1}^{k}}{m^{k}}\cdot\|\partial_{i}^{k}g\|_{L^{\infty}([0,\pi]^{d})}\leq\frac{dC_{1}^{k}}{m^{k}}\cdot\max_{|{\boldsymbol{\alpha}}|=k}\|\partial^{\boldsymbol{\alpha}}g\|_{L^{\infty}([0,\pi]^{d})}.$
For the entry-wise bound on the Fourier coefficients, note that
$Q({\boldsymbol{\theta}})=\frac{1}{\pi^{d}}\int_{[0,\pi]^{d}}\prod_{i=1}^{d}J_{1}(y_{i})G(\theta_{1}-y_{1},\dots,\theta_{d}-y_{d})\mathrm{d}{\mathbf{y}}.$
And we can compute the coefficients
$\tilde{c}_{K}$ $=\frac{2^{\mathrm{nnz}(K)}}{(2\pi)^{d}}\int_{[-\pi,\pi]^{d}}Q({\boldsymbol{\theta}})\prod_{i=1}^{d}\cos(k_{i}\theta_{i})\mathrm{d}{\boldsymbol{\theta}}$
$=\frac{2^{\mathrm{nnz}(K)}}{(2\pi)^{d}}\int_{[-\pi,\pi]^{d}}\frac{1}{\pi^{d}}\int_{[0,\pi]^{d}}\prod_{i=1}^{d}J_{1}(y_{i})G(\theta_{1}-y_{1},\dots,\theta_{d}-y_{d})\mathrm{d}{\mathbf{y}}\prod_{i=1}^{d}\cos(k_{i}\theta_{i})\mathrm{d}{\boldsymbol{\theta}}$
$=\frac{1}{\pi^{d}}\int_{[0,\pi]^{d}}\prod_{i=1}^{d}J_{1}(y_{i})\;\frac{2^{\mathrm{nnz}(K)}}{(2\pi)^{d}}\int_{[-\pi,\pi]^{d}}G(\theta_{1}-y_{1},\dots,\theta_{d}-y_{d})\prod_{i=1}^{d}\cos(k_{i}\theta_{i})\mathrm{d}{\boldsymbol{\theta}}\;\mathrm{d}{\mathbf{y}}$
$=\frac{1}{\pi^{d}}\int_{[0,\pi]^{d}}\prod_{i=1}^{d}J_{1}(y_{i})\;\frac{2^{\mathrm{nnz}(K)}}{(2\pi)^{d}}\int_{[-\pi,\pi]^{d}}G(u_{1},\dots,u_{d})\prod_{i=1}^{d}\cos(k_{i}(u_{i}+y_{i}))\mathrm{d}{\mathbf{u}}\;\mathrm{d}{\mathbf{y}}$
$\overset{\text{(a)}}{=}\frac{1}{\pi^{d}}\int_{[0,\pi]^{d}}\prod_{i=1}^{d}J_{1}(y_{i})\cos(k_{i}y_{i})\mathrm{d}{\mathbf{y}}\cdot\frac{2^{\mathrm{nnz}(K)}}{(2\pi)^{d}}\int_{[-\pi,\pi]^{d}}G(u_{1},\dots,u_{d})\prod_{i=1}^{d}\cos(k_{i}u_{i})\mathrm{d}{\mathbf{u}}$
$=\frac{1}{\pi^{d}}\int_{[0,\pi]^{d}}\prod_{i=1}^{d}J_{1}(y_{i})\cos(k_{i}y_{i})\mathrm{d}{\mathbf{y}}\cdot c_{K}$
$=c_{K}\prod_{i=1}^{d}\Big(\frac{1}{\pi}\int_{0}^{\pi}J_{1}(y_{i})\cos(k_{i}y_{i})\mathrm{d}y_{i}\Big).$
Here in the step (a), we expand $\cos(k_{i}(u_{i}+y_{i}))=\cos(k_{i}u_{i})\cos(k_{i}y_{i})-\sin(k_{i}u_{i})\sin(k_{i}y_{i})$, and we drop the sine term as $G$ is an even function and its integral against $\sin(k_{i}y_{i})$ would vanish. And then $|\tilde{c}_{K}|\leq|c_{K}|$ follows since
$\Big|\frac{1}{\pi}\int_{0}^{\pi}J_{1}(y_{i})\cos(k_{i}y_{i})\mathrm{d}y_{i}\Big|\leq\frac{1}{\pi}\int_{0}^{\pi}J_{1}(y_{i})\mathrm{d}y_{i}=1.$
Finally, we will complete the proof by proving the properties for $f({\mathbf{x}})$. By changes of variables we have $f({\mathbf{x}})=\sum_{K}c_{K}\overline{T}_{K}({\mathbf{x}})$, and we define
$\tilde{f}({\mathbf{x}})=\sum_{K\in\{0,\dots,m\}^{d}}\tilde{c}_{K}\overline{T}_{K}({\mathbf{x}}).$
Then $|\tilde{c}_{K}|\leq|c_{K}|$ still holds. Moreover, we have
$\|f-\tilde{f}\|_{L^{\infty}([-1,1]^{d})}=\|g-\tilde{g}\|_{L^{\infty}([0,\pi]^{d})}\leq\frac{dC_{1}^{k}}{m^{k}}\cdot\max_{|{\boldsymbol{\alpha}}|=k}\|\partial^{\boldsymbol{\alpha}}g\|_{L^{\infty}([0,\pi]^{d})}.$
In the proof of Proposition 1, we have shown that
$\|\partial^{\boldsymbol{\alpha}}g\|_{L^{\infty}([0,\pi]^{d})}\leq|{\boldsymbol{\alpha}}|!\cdot\left(1+\frac{k}{d}\right)^{d}\,\max_{|{\boldsymbol{\beta}}|\leq k}\|\partial_{x}^{\boldsymbol{\beta}}f\|_{L^{\infty}([-1,1]^{d})}\leq k!\cdot e^{k}.$
Taking maximum over $\boldsymbol{\alpha}$} concludes the proof. ∎
Appendix E Estimation for $S$.
Lemma 13.
Let
$S\coloneqq\sum_{K\in\{0,\dots,m\}^{d}\setminus\{\mathbf{0}\}}\frac{1}{\|K\|_{2}^{k}}.$
Then the following holds:
$S\;\leq\;\begin{cases}d\,2^{d-1}\!\left(1+\dfrac{m^{\,d-k}-1}{d-k}\right),&\text{if }k<d,\\ d\,2^{d-1}\,(1+\log m),&\text{if }k=d,\\ d\,2^{d-1}\!\left(1+\dfrac{1}{k-d}\right),&\text{if }k>d.\end{cases}$
Proof.
For $j=1,\dots,m$, define the shell
$\mathcal{A}_{j}:=\bigl\{K\in\{0,\ldots,m\}^{d}:\ \|K\|_{\infty}=j\bigr\},\qquad\|K\|_{\infty}:=\max_{1\leq i\leq d}|K_{i}|.$
Then
$S=\sum_{j=1}^{m}\ \sum_{K\in\mathcal{A}_{j}}\frac{1}{\|K\|_{2}^{\,k}}.$
For $K\in\mathcal{A}_{j}$ we have $j\leq\|K\|_{2}\leq\sqrt{d}\,j$, hence
$\frac{1}{\|K\|_{2}^{\,k}}\leq\frac{1}{j^{k}}\qquad\Rightarrow\qquad S\leq\sum_{j=1}^{m}\frac{|\mathcal{A}_{j}|}{j^{k}}.$
Points with $\|K\|_{\infty}\leq j$ form a $(j+1)^{d}$ grid, hence $|\mathcal{A}_{j}|=(j+1)^{d}-j^{d}$. By the mean value theorem applied to $x\mapsto x^{d}$, there exists $\xi\in(j,j+1)$ such that
$(j+1)^{d}-j^{d}=d\,\xi^{d-1}\leq d\,(j+1)^{d-1}\leq d\,2^{d-1}j^{d-1}\qquad\text{when }j\geq 1.$
Therefore
$S\;\leq\;d\,2^{\,d-1}\sum_{j=1}^{m}j^{\,d-1-k}.$ (28)
Let $\alpha:=d-1-k$. Using the integral bound:
  • If $k<d$ (i.e. $\alpha>-1$), then
    $\sum_{j=1}^{m}j^{\alpha}\;\leq\;1+\int_{1}^{m}x^{\alpha}\,dx\;=\;1+\frac{m^{\alpha+1}-1}{\alpha+1}\;=\;1+\frac{m^{\,d-k}-1}{d-k}.$
    Substituting into (28) gives
    $S\;\leq\;d\,2^{\,d-1}\!\left(1+\frac{m^{\,d-k}-1}{d-k}\right).$
  • If $k=d$ (i.e. $\alpha=-1$), then
    $\sum_{j=1}^{m}\frac{1}{j}\;\leq\;1+\int_{1}^{m}\frac{dx}{x}\;=\;1+\log m,$
    and (28) yields $S\leq d2^{\,d-1}\,(1+\log m)$.
  • If $k>d$ (i.e. $\alpha<-1$), then
    $\sum_{j=1}^{\infty}j^{\alpha}\;\leq\;1+\int_{1}^{\infty}x^{\alpha}\,dx\;=\;1+\frac{1}{k-d},$
    hence $\sum_{j=1}^{m}j^{\alpha}\leq 1+\frac{1}{k-d}$ for all $m\geq 1$ and
    $S\leq d2^{d-1}\!\left(1+\frac{1}{k-d}\right).$
This finishes the proof for all three cases. ∎
Appendix F Computational complexity of Algorithm 1
The running time of Algorithm 1 is dominated by the optimization step, which solves a constrained least-squares problem involving $m^{kd}$ variables and $m^{d}-1$ moment-matching constraints. With the choice of $m$ specified in Section 4.2, this leads to an overall time complexity of
$O\left((\varepsilon n)^{(k+2)\min\{k,d\}}\right).$
We refer to [28] for background on the computational complexity of large-scale least-squares problems. Finally, we note that the optimization problem exhibits additional tensor structure that could potentially be exploited to further reduce the computational cost; exploring such improvements, however, is beyond the scope of this work.